OpenStudy (anonymous):

A 5-card hand is dealt from a deck of 52 cards. What is the probability that 3 cards are queens and two are kings?

OpenStudy (amistre64):

qqqkk has 5!/(3! 2!) ways of getting pulled 4P3 queens, 4P2 kings = 4.3.2.4.3; out of 52P5 outcomes

OpenStudy (amistre64):

so i see it as:\[10*\frac{4.5.3.4.3}{52.51.50.49.48}\]simplify as needed

OpenStudy (anonymous):

I'm a little confused on how you did it. Also the answers around 0.00000923

OpenStudy (amistre64):

got a typo in th enumerator ...

OpenStudy (amistre64):

i added up all the possible ways to pull 3 queens and 2 kings: qqqkk1 qqkqk2 qkqqk3 kqqqk4 qqkkq5 qkqkq6 kqqkq7 qkkqq8 kqkqq9 kkqqq10 the probability of something happening is the sum of all the ways it can happen as we pull the cards we have 4queens to pick from, then 3, then 2 we have 4 kings to pick from , then 3 giving us the respective probabilities for a single permuations

OpenStudy (amistre64):

4P3 = 4.3.2 4P2 = 4.3 10 ways to pull them, and 52 pick 5 different ways to pull 5 cards out of 52

OpenStudy (amistre64):

\[10*\frac{4.3.2.4.3}{52.51.50.49.48}\] \[\frac{4.3.2.4.3}{52.51.5.49.48}\] \[\frac{3.2.4.3}{52.51.5.49.12}\] \[\frac{3.2}{52.51.5.49}\] \[\frac{3}{26.51.5.49}\] \[\frac{1}{26.17.5.49}\]

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=1%2F%2826*17*5*49%29 9.2344630... × 10^-6

OpenStudy (anonymous):

Thank you!

OpenStudy (amistre64):

youre welcomme